OpenStudy (anonymous):

Find a function f and a number a such that 2+∫ a to x f(t)/t^4 dt = 6x^−1 f(x) = a =

myininaya (myininaya):

\[2+\int\limits_a^x \frac{f(t)}{t^4} dt=6x^{-1}\] i think I would try differentiating both sides first

myininaya (myininaya):

you will first be able to find f(x) this way or f(t) if you replace the x's with t's later then you will plug back and and solve the equation for a

myininaya (myininaya):

if you want you can let me know what you have after differentiating both sides

OpenStudy (anonymous):

well the derivative for the right side is -6x^-2. i'm not sure how to differentiate the left side :(

myininaya (myininaya):

fundamental theorem of calculus you did it last problem

myininaya (myininaya):

derivative of 2 is 0 of course now look at the integral and differentiate that part

OpenStudy (anonymous):

f(x)/ x^4 or is there something more?

myininaya (myininaya):

\[\frac{d}{dx} \int\limits _a^xg(t) dt=g(x)\] this one yes! !!! that is it

myininaya (myininaya):

solve for f(x) now

myininaya (myininaya):

you have \[\frac{f(x)}{x^4}=-6x^{-2}\]

OpenStudy (anonymous):

-6x^2

myininaya (myininaya):

beautiful

OpenStudy (anonymous):

:)

myininaya (myininaya):

now plug f(t)=-6t^2 in and solve for a \[2+\int\limits\limits_a^x \frac{-6t^2}{t^4} dt=6x^{-1}\]

myininaya (myininaya):

let me know if you need more help or if you are confused

OpenStudy (anonymous):

yeah i need help. can you show me the process?

myininaya (myininaya):

you can simplify the integrand then use power rule to integrate then plug in limits

myininaya (myininaya):

\[-6 \frac{t^2}{t^4}=-6t^{-2}\] is integrand after simplification

myininaya (myininaya):

integrate this integrand by use of power rule (power rule for integrals that is since this is an integral :p)

OpenStudy (anonymous):

i'm kind of new to integrals, can you show mw what to do?

myininaya (myininaya):

\[2+\int\limits\limits\limits_a^x \frac{-6t^2}{t^4} dt=6x^{-1} \\ 2+\int\limits_a^x -6t^{-2} dt=6x^{-1} \\ \text{ you are evaluating the integral right now }\] recall \[n \neq -1 \\ \int\limits_{a}^{b}x^{n} dx=\frac{x^{n+1}}{n+1}|_a^b=\frac{a^{n+1}}{n+1}-\frac{b^{n+1}}{n+1}\]

myininaya (myininaya):

oops I pluggin in a first didn't mean to do that

myininaya (myininaya):

\[n \neq -1 \\ \int\limits\limits_{a}^{b}x^{n} dx=\frac{x^{n+1}}{n+1}|_a^b=\frac{b^{n+1}}{n+1}-\frac{a^{n+1}}{n+1}\]

myininaya (myininaya):

@lawls you okay?

OpenStudy (anonymous):

yeah i just went somewhere sorry to keep you waiting

myininaya (myininaya):

here is a big hint if you return: you might end up with something like this after doing the evaluation of the definite integral: \[2+g(x)-g(a)=g(x) \]

myininaya (myininaya):

oops I was typing that before I noticed you came back lol

OpenStudy (anonymous):

lol ok let me try to read what you wrote

OpenStudy (anonymous):

i got [a^-1 / -1] - [x^-1/-1]

myininaya (myininaya):

hmm do you mean this: \[2+-6 \int\limits_a^x t^2 =6x^{-1} \\ 2+6 \int\limits_x^a t^2 =6x^{-1} \\ 2+6[-t^{-1}]_x^a=6x^{-1} \\ 2+6(-a^{-1}+x^{-1})=6x^{-1}\] if so yes that is the right track

myininaya (myininaya):

you can distribute the 6 there

myininaya (myininaya):

and cancel out the functions of x's just like we have 2+g(x)-g(a)=g(x) we can cancel the g(x) by subtracting g(x) on both sides 2-g(a)=0

myininaya (myininaya):

so you will have an equation just in terms of a

OpenStudy (anonymous):

ok a =3

myininaya (myininaya):

yeah that is what I have too :) gj

OpenStudy (anonymous):

thank you so much!!

myininaya (myininaya):

np